On April 30, 2013, McKensie Products purchased machinery for $66,000. The useful life of this machinery is estimated at 5 years, with a $16,000 residual value. The company uses the double-declining-balance method. Depreciation expense for the fiscal year ending on December 31, 2014 will be:

  1. $15,840
  2. $18,400
  3. $21,067
  4. $19,360
Answer #1

Answer :$19,360

Double Decline Balancing Method

Depreciation Rate for Year = (1/Estimated)*2 = (1/5)*2 = 40%

Cost of the Machinery = $66,000

Deprecition as on December 31st 2013 = $66,000*40%*8/12 = $17,600

Depreciation As on December 31st 2014 =($66,000-$17,600)*40% = $19,360 (Answer)

Note : For Any Further Clarification Plases Use the Comment Box.Thank You Very much